# Token-Budget-Aware LLM Reasoning

### v3

@@ Abstract
Reasoning is critical for large language models (LLMs) to excel in a wide range of tasks. While methods like Chain-of-Thought (CoT) reasoning enhance LLM performance by decomposing problems into intermediate steps, they also incur significant overhead in token usage, leading to increased costs. We find that the reasoning process of current LLMs is unnecessarily lengthy and it can be compressed by including a reasonable token budget in the prompt, but the choice of token budget plays a crucial role in the actual compression effectiveness. We then propose a token-budget-aware LLM reasoning framework, which dynamically estimates token budgets for different problems based on reasoning complexity and uses the estimated token budgets to guide the reasoning process. Experiments show that our method effectively reduces token costs in CoT reasoning with only a slight performance reduction, offering a practical solution to balance efficiency and accuracy in LLM reasoning. Code: https://github.com/GeniusHTX/TALE.
@@
footnotetext: Pre-print with preliminary results, work in progress.
@@ 1 Introduction
Reasoning plays a crucial role in enabling large language models (LLM) to perform effectively across a wide range of tasks Zhou et al. (2022); Hao et al. (2023, 2024a). A variety of methods have been proposed to enhance the reasoning capabilities of large language models Suzgun et al. (2022); Wang et al. (2023); Feng et al. (2023); Yao et al. (2024a); Xie et al. (2024). Among these, Chain-of-Thought (CoT) Wei et al. (2022) is the most representative and widely adopted approach. It enhances the reliability of the model’s answers by guiding large language models with the prompt "Let’s think step by step", encouraging them to decompose the problem into intermediate steps and solve each before arriving at the final answer. Figure 1a and Figure 1b illustrate an intuitive example. Observe that without CoT, the LLM produces incorrect answers to the question. With a CoT-enhanced prompt, the LLM systematically breaks the question into multiple steps and reasons through each step sequentially. By addressing each step incrementally, the LLM eventually arrives at the correct answer.
@@
Figure 1: Examples of different problem solving paradigms. The reasoning processes are highlighted.
Although reasoning enhancement approaches such as CoT impressively improve LLM performance, they produce substantial additional overhead, specifically in the form of the increased number of tokens produced Wei et al. (2022); Feng et al. (2023); Yao et al. (2024a). As shown in Figure 1b, the answer to prompt with CoT has notably higher token costs due to the detailed intermediate reasoning steps included in the output. Such high token costs can lead to significant expenses, including increased computational resource usage and longer running times during the LLM inference phase, ultimately resulting in significant additional monetary and energy costs.
@@
Based on the above observations and analysis, we designed a prototype for token-budget-aware reasoning in large language models (LLMs). Our approach leverages the token budget to guide the reasoning process, dynamically allocating different token budgets to problems based on an estimation of their reasoning complexity. We call our method TALE (Token-Budget-Aware LLM rEasoning). For a given problem and a specific LLM, TALE first estimates an appropriate token budget and then uses it to guide the reasoning process. We discuss different implementations of TALE in Section 5. Experiment results show that TALE significantly reduces token costs in LLM chain-of-thought (CoT) reasoning while largely maintaining the correctness of the answers. On average, TALE achieves a 68.64% reduction in token usage while maintaining accuracy with less than 5% decrease.
@@ 4 Searching Optimal Token Budget
Vanilla Method for Optimal Budget Search. An intuitive method is finding the minimal needed tokens as the budget, ensuring that the LLM can still produce correct and accurate responses within this constraint. To find the minimal token budget required for each question, we utilize a binary search-based minimal budget search algorithm. Algorithm 1 showcases the details. Before initiating the search process, we first apply the vanilla CoT to generate an answer for each question, as illustrated in Figure 1b. The number of tokens in the resulting answer is then calculated and designated as the right boundary for search, denoted by $right$. The function isFeasible is used to determine the feasibility of a budget. A budget is considered feasible here if the CoT prompt with that budget preserves the correctness of the answer. Given the feasibility function, large language model $\mathcal{M}$, question ${\bm{x}}$ and label $y$as the input, Algorithm 1 first calculates the right boundary of search (line 2). With 00 as the left boundary, the current possible budget $\beta$ is computed as the midpoint of 00 and $right$ (line 3). We use $\beta_{0}$ to record the previously searched budget (line 4). While the current $\beta$ is feasible, the algorithm updates $\beta$ by recalculating the midpoint (line 7) and adjusts the search bounds accordingly to narrow the range (line 9). Once the loop ends, the final budget $\beta$ is returned as the searched result (line 12). Algorithm 1 is designed to find the minimal budget efficiently. However, we observe that the minimal budget required to produce a correct answer is not necessarily the optimal budget. When the budget is unreasonably small, the actual token cost often exceeds that of cases where a larger budget is used.
@@
Observation of Token Elasticity. During our minimal budget search process, we observe a “token elasticity” phenomenon as we approach the minimal budget. Specifically, as Algorithm 1 progresses, we aim to identify the minimal budget that still ensures the answer’s correctness. However, we find that if the budget is reduced beyond a certain range, the token cost increases, indicating that further reductions in the budget lead to increasing token consumption. Figure 2 showcases the evidence. The x-axis represents the iterations of the budget binary search, with the budget values decreasing progressively. The y-axis in Figure 2b and Figure 2d show the corresponding token costs at each budget search iteration. We can observe that when the searched budget crosses a range for reasonable budgets, the token costs begin to gradually increase. Figure 1c also shows an example. As observed, when a small token budget (e.g., 10 tokens) is used, the real token cost is significantly higher compared to scenarios where a reasonable token budget is allocated (i.e., Figure 1d).
@@
Token Elasticity based Optimal Budget Search. The token elasticity observation shows that while a minimal budget may keep the correctness of the answer, it does not necessarily minimize the token cost. Figure 1c and Figure 1d illustrate an intuitive example. To address this, we enhance Algorithm 1 by incorporating a greedy search strategy aimed at finding the optimal budget that simultaneously minimizes token cost and preserves answer correctness. Specifically, we introduce an additional constraint to the isFeasible condition. Beyond ensuring correctness, the updated budget must result in a lower token cost compared to the previously searched budget. Algorithm 2 outlines the feasibility function employed during the search process. Initially, the actual token cost is computed for both the current and previously evaluated budgets (line 2). Next, feasibility is assessed based on two criteria: the correctness of the answer and a reduction in token cost following a greedy strategy (line 3). The search process is terminated if either condition fails.
Figure 4: The workflow of TALE. Given a question, TALE first estimates the token budget using a budget estimator. It then crafts a token-budget-aware prompt by combining the question with the estimated budget. Finally, the prompt is input to the LLM to generate the answer as the final output.
@@ 5.1 Overview
Based on the above observations and analysis, we designed our method TALE for token-budget-aware reasoning in LLMs. Figure 4 provides an overview of TALE’s workflow. TALE aims to elaborate a token-budget-aware prompt that achieves performance comparable to vanilla CoT while reducing token costs. To strike this balance, TALE follows a two-phase approach: budget estimation and prompt construction. The observation of token elasticity, as discussed in Section 4, suggests that only an appropriate budget located in a reasonable range can effectively reduce token costs and preserve LLM performance simultaneously. The optimal budget searched by Algorithm 1 and Algorithm 2 is located in such reasonable budget range and achieves satisfying trade-off between token costs and LLM performance. In this case, TALE first estimates a reasonable token budget that is close to the searched optimal budget for the given question. Using the estimated token budget, TALE then crafts a token-budget-aware prompt and then feeds it into LLM to generate the final answer. Figure 7 illustrates an example of TALE.
5.2 Budget Estimation
To estimate an appropriate budget within the reasonable budget range, two possible solutions are taken into consideration: zero-shot-based mechanism and budget regression. We also discuss an approach that internalizes the budget awareness of LLM by fine-tuning it. All approaches focus on developing an estimator that effectively approximates the optimal budget.
Zero-shot Estimator. For zero-shot mechanism to predict the reasonable budget, TALE leverages the reasoning LLM itself as an estimator. Before querying the LLM with a question, TALE first prompts the LLM to estimate the number of output tokens needed to answer the question. Figure 5 illustrates the budget estimation prompt. The key intuition behind this is the human-like thinking paradigm. When presented with a mathematical question, although it may take humans a few minutes to calculate the answer, they can typically estimate the time or effort required to solve it after just briefly reviewing the question. For instance, when presented with a question from primary school arithmetic and another from college-level calculus, a human may not immediately provide the answers. Still, it is easy to infer that the former can be solved in seconds, while the latter requires a significantly longer time, even with only a brief glance. RQ2 in Section 6 showcases the performance of budget estimation. Observe that a large amount of the estimated budgets are around the optimal searched budget and achieve competitive performance.
Figure 5: The prompt for zero-shot estimator.
Regression Estimator. For regression-based estimator, we aim to train/fine-tune another LLM $f(\theta)$ to serve as the estimator, such that $f(\theta)$ estimates the optimal token budget given a specific LLM and a particular question. Given $\mathcal{D}$ = $\{({\bm{p}},{\bm{x}}_{i},\beta_{i}^{*})\}_{i=1}^{N}$, ${\bm{p}}$ is the instruction prompt, ${\bm{x}}_{i}$ is a question, $\beta_{i}^{*}$ is our searched optimal budget (searched by Algorithm 1 and Algorithm 2) for ${\bm{x}}$ and $N$ is the dataset size. We assign the instruction prompt ${\bm{p}}$ as “Estimate the token budget for the following question”. Next, we initialize $f(\theta)$ using a pre-trained LLM, such as LLaMA 3-8B AI@Meta (2024). Then, we craft the target output $y_{i}$ for ${\bm{x}}_{i}$ using $\beta_{i}^{*}$. For example, given a $\beta_{0}^{*}$ as 14, the corresponding target output $y_{0}$ is “The token budget is 14”. We aim to make the model output as close as possible to the target output $y_{i}$ by minimizing loss function, which can be formalized as:
$\theta^{*}=\arg\min_{\theta}\sum_{i=1}^{N}\mathcal{L}\left(f_{\theta}({\bm{p}}_{i},{\bm{x}}_{i}),y_{i}\right)$ (1)
where $\theta$ denotes the model parameters and $\theta^{*}$ the optimized parameters. The loss function is defined as follows:
$\mathcal{L}=-\sum_{i=1}^{N}\log\mathbb{P}(y_{i}\mid{\bm{p}}_{i},{\bm{x}}_{i};\theta)$ (2)
Through Equation 2, model parameters $\theta$ are optimized to maximize the probability $P(y_{i}\mid{\bm{p}}_{i},{\bm{x}}_{i};\theta)$ of the target output $y_{i}$ given the input ${\bm{x}}_{i}$ and instruction prompt ${\bm{p}}$ across all $N$ training samples.
@@ 5.2 Budget Estimation
Token-Budget Awareness Internalization. To obtain an LLM with token awareness, we fine-tune the LLM to internalize the budget estimation into the inference process and produce token-efficient reasoning responses. Specifically, we fine-tune the LLM $\mathcal{M}(\theta)$ so that it generates token-budget-aware answers. This process is divided into two key stages: target output generation and LLM fine-tuning. In the target output generation stage, we craft the target output $y_{i}$ by prompting $\mathcal{M}(\theta)$ with a Chain-of-Thought (CoT) prompt that incorporates our searched optimal token budget. The prompt is formatted as follows:
@@
where $\beta_{i}^{*}$ is the searched optimal budget for the given question ${\bm{x}}_{i}$ (see search process in Algorithm 1 and Algorithm 2). Figure 1d illustrates an example. The resulting LLM output, constrained by the token budget specified in the prompt, is taken as the crafted target output $y_{i}$. In the LLM fine-tuning stage, we train the LLM $\mathcal{M}(\theta)$ using the crafted target outputs from the first stage. The instruction prompt for fine-tuning is standardized as follows:
‘‘Let’s think step by step:’’
The training sample for fine-tuning consists of the given question ${\bm{x}}_{i}$ paired with the crafted target output $y_{i}$. The training objective aligns with the loss function described in Equation 2 where $({\bm{x}},y_{i})$ indicates the training sample and ${\bm{p}}_{i}$ means the above instruction prompt. During fine-tuning, the LLM is encouraged to internalize the token budget constraint and prefer a compact reasoning process, following the target outputs generated in the first stage. This two-stage process ensures that the LLM produces concise yet correct responses, effectively balancing reasoning quality with token efficiency during inference.
6 Evaluation
In this section, we provide the preliminary results of the zero-shot estimator version of our method TALE. This project is ongoing and we will update more results soon. Three state-of-the-art LLMs (i.e., GPT-4o OpenAI (2024b), GPT-4o-mini OpenAI (2024a), Yi-lightning Wake et al. (2024)) are involved in the experiments. Our evaluation is surrounding around the following research questions(RQs):
RQ1. How effective is TALE in reducing token costs while maintaining LLM performance?
RQ2. How effective is TALE to estimate the token budget given a question?
RQ3. How general is TALE across different state-of-the-art LLMs?
6.1 Experiment Setup
Metrics. The target of TALE is to balance the LLM performance and extra redundant token costs. Specifically, TALE seeks to minimize token consumption while maintaining comparable LLM performance. To evaluate the LLM performance, three most challenging mathematical datasets are taken into consideration: GSM8K Cobbe et al. (2021), GSM8K-Zero Chiang and Lee (2024), and MathBench Liu et al. (2024). GSM8K-Zero, derived from the GSM8K dataset, specifically targets the analysis of over-reasoning and redundancy in LLM-generated outputs. In short, GSM8K-Zero is designed so that the answers are embedded within the questions themselves. LLMs can easily generate correct responses without complicated additional reasoning or redundant calculations.
Number of Output Tokens. We evaluate the token costs by calculating the average output token consumption for each specific task. The output token costs are measured as follows: Number of Output $\textit{Number of Output Tokens}=\frac{1}{N}\sum_{i=1}^{N}{\mathbb{T}(\mathcal{M}({\bm{x}}_{i}))}$, where ${\bm{x}}_{i}$ represents the given question, and $\mathbb{T}$ is a function that measures the number of tokens. Intuitively, the more output tokens, the higher the costs incurred by $\mathcal{M}$. To evaluate costs more precisely, we calculate the average expense per sample for querying the LLM. The total token expense includes both input and output tokens used during the query process.
6.2 RQ1. Effectiveness of TALE.
Table 2 presents a comparison of TALE and other prompt engineering methods, including “Directly Answering” and “Vanilla CoT”, in their effectiveness for seven different datasets. The results are evaluated regarding ACC, Number of Output Tokens (Output Tokens for short), and Expense. A well-designed prompt engineering approach should induce the LLM to generate a correct response with as few tokens as possible. TALE consistently demonstrates significant improvements in efficiency while maintaining competitive accuracy. Directly answering achieves the lowest output tokens (14.57 on average) and expenses (25.37 on average) but has the lowest accuracy (52.31% on average). Vanilla CoT achieves the highest accuracy (83.75% on average) but incurs a significant token cost (461.25 on average) and expenses (289.78 on average). TALE demonstrates a balanced trade-off between performance and cost. It achieves competitive accuracy (81.03%) while reducing token costs (32% of Vanilla CoT) and expenses (41% of Vanilla CoT) significantly. For accuracy, notably, on GSM8K, TALE even improves accuracy to 84.46%, surpassing Vanilla CoT, demonstrating its ability to adapt well to complex reasoning tasks while remaining efficient. For output tokens on GSM8K-Zero, TALE achieves an impressive reduction in output token costs from 252.96 to 22.67 while maintaining high accuracy (98.72%), showcasing its capability to optimize reasoning efficiency in such tasks. For expenses, TALE demonstrates its cost-effectiveness, reducing expenses from 78.58 to 18.62 while achieving reasonable accuracy (73.67% vs 75.00%) on MathBench-Arithmetic. TALE demonstrates that incorporating token-budget awareness allows for a significant reduction in token costs and monetary expenses without a major compromise in accuracy. TALE reduces output token costs by 68.64% on average, making it a more efficient solution for budget-constrained reasoning tasks while retaining competitive performance. These results highlight TALE’s generalizability across tasks with varying complexity, demonstrating its potential to scale in real-world scenarios while managing computational and financial resources effectively.
@@ 6.1 Experiment Setup
Table 2: Comparison of TALE (Zero-shot Estimator Version) and other prompt engineering methods. “Directly Answering” means prompting LLM without any reasoning process. “Vanilla CoT” means the vanilla CoT prompting with budget. The model used in our evaluation is GPT-4o-mini OpenAI (2024a). Observe that TALE achieves an average accuracy (ACC) of 80.22%, with an average output token cost of 138.53 and an average expense of 118.46. TALE reduces output token costs by 67%, lowers expenses by 59%, and maintains competitive performance compared to the vanilla CoT approach. ACC $\uparrow$, Output Tokens $\downarrow$, Expense ($10^{-5}\$$ / sample) $\downarrow$.
Dataset Directly Answering Vanilla CoT TALE (Ours)
@@
Table 3: The generalization of TALE (Zero-shot Estimator Version) across different LLMs. Yi-lightning Wake et al. (2024), GPT-4o-mini OpenAI (2024a) and GPT-4o OpenAI (2024b) are taken into consideration. We conduct the evaluation on MathBench-College. ACC $\uparrow$, Output Tokens $\downarrow$, Expense ($10^{-5}\$$ / sample) $\downarrow$.
LLM Directly Answering Vanilla CoT TALE (Ours)

### v4

@@ Abstract
Reasoning is critical for large language models (LLMs) to excel in a wide range of tasks. While methods like Chain-of-Thought (CoT) reasoning and enhance LLM performance by decomposing problems into intermediate steps, they also incur significant overhead in token usage, leading to increased costs. We find that the reasoning process of current LLMs is unnecessarily lengthy and it can be compressed by including a reasonable token budget in the prompt, but the choice of token budget plays a crucial role in the actual compression effectiveness. We then propose a token-budget-aware LLM reasoning framework that dynamically adjusts the number of reasoning tokens based on the reasoning complexity of each problem. Experiments show that our method effectively reduces token costs in CoT reasoning with only a slight performance reduction, offering a practical solution to balance efficiency and accuracy in LLM reasoning. Code: https://github.com/GeniusHTX/TALE.
@@ 1 Introduction
Reasoning plays a crucial role in enabling large language models (LLM) to perform effectively across a wide range of tasks Zhou et al. (2022); Hao et al. (2023, 2024a). A variety of methods have been proposed to enhance the reasoning capabilities of large language models Suzgun et al. (2022); Wang et al. (2023); Feng et al. (2023); Xie et al. (2024). Among these, Chain-of-Thought (CoT) Wei et al. (2022) is the most representative and widely adopted approach. It enhances the reliability of the model’s answers by guiding large language models with the prompt “Let’s think step by step”, encouraging them to decompose the problem into intermediate steps and solve each before arriving at the final answer. Figure 1a and Figure 1b illustrate an intuitive example. Observe that without CoT, the LLM produces incorrect answers to the question. With a CoT-enhanced prompt, the LLM systematically breaks the question into multiple steps and reasons through each step sequentially. By addressing each step incrementally, the LLM eventually arrives at the correct answer. Recent reasoning models, such as OpenAI O1 OpenAI (2024c) and DeepSeek R1 Guo et al. (2025), integrate CoT into their design. Notably, these models can perform CoT reasoning even without explicit prompting.
@@
Figure 1: Examples of different problem solving paradigms. The reasoning processes are highlighted. We conduct this experiment on GPT-4o-mini.
Although reasoning enhancement approaches such as CoT impressively improve LLM performance, they produce substantial additional overhead, specifically in the form of the increased number of tokens produced Wei et al. (2022); Feng et al. (2023); Yao et al. (2024a). As shown in Figure 1b, the answer to prompt with CoT has notably higher token costs due to the detailed intermediate reasoning steps included in the output. Such high token costs can lead to significant expenses, including increased computational resource usage and longer running times during the LLM inference, ultimately resulting in significant additional monetary and energy costs.
@@
Based on the above observations and analysis, we propose a token-budget-aware LLM reasoning framework that dynamically adjusts the number of reasoning tokens based on the reasoning complexity of each problem. We call our method TALE (Token-Budget-Aware LLM rEasoning), which includes two implementations: token budget estimation and prompting (TALE-EP) and token budget awareness internalization via post-training (TALE-PT). TALE-EP estimates a reasonable token budget for each problem using zero-shot prompting and incorporates it into the reasoning process, while TALE-PT internalizes token-budget awareness through post-training, enabling the LLM to generate more token-efficient responses without explicit token constraints in the prompt. We discuss both implementations in Section 5. Experiment results show that TALE significantly reduces token costs in LLM chain-of-thought (CoT) reasoning while largely maintaining answer correctness. On average, TALE-EP achieves a 67% reduction in token usage while maintaining accuracy with less than a 3% decrease. TALE-PT cuts token usage by around 50% compared to Vanilla CoT and achieves competitive performance.
@@ 4 Searching Optimal Token Budget
Vanilla Method for Optimal Budget Search. An intuitive method is finding the minimal needed tokens as the budget, ensuring that the LLM can still produce correct and accurate responses within this constraint. To find the minimal token budget required for each question, we utilize a binary search-based minimal budget search algorithm. Before initiating the search process, we first apply the vanilla CoT to generate an answer for each question, as illustrated in Figure 1b. The number of tokens in the resulting answer is then calculated and designated as the right boundary for search, denoted by $right$. The function isFeasible is used to determine the feasibility of a budget. A budget is considered feasible here if the CoT prompt with that budget preserves the correctness of the answer. Algorithm 1 showcases the details. Given the feasibility function, large language model $\mathcal{M}$, question ${\bm{x}}$ and label $y$ as the input, Algorithm 1 first calculates the right boundary of search (line 2). With 00 as the left boundary, the current possible budget $\beta$ is computed as the midpoint of 00 and $right$ (line 3). We use $\beta_{0}$ to record the previously searched budget (line 4). While the current $\beta$ is feasible, the algorithm updates $\beta$ by recalculating the midpoint (line 7) and adjusts the search bounds accordingly to narrow the range (line 9). Once the loop ends, the final budget $\beta$ is returned as the searched result (line 12). Algorithm 1 is designed to find the minimal budget efficiently. However, we observe that the minimal budget required to produce a correct answer is not necessarily the optimal budget. When the budget is unreasonably small, the actual token cost often exceeds that of cases where a larger budget is used.
@@
Observation of Token Elasticity. During our minimal budget search process, we observe a “token elasticity” phenomenon as we approach the minimal budget. Specifically, as Algorithm 1 progresses, we aim to identify the minimal budget that still ensures the answer’s correctness. However, we find that if the budget is reduced beyond a certain range, the token cost increases, indicating that further reductions in the budget lead to increasing token consumption. Figure 2 showcases the evidence. The x-axis represents the iterations of the budget binary search, with the budget values decreasing progressively. The y-axis in Figure 2b and Figure 2d show the corresponding token costs at each budget search iteration. We can observe that when the searched budget crosses a range for reasonable budgets, the token costs begin to increase gradually. Figure 1c also shows an example. As observed, when a small token budget (e.g., 10 tokens) is used, the real token cost is significantly higher compared to scenarios where a reasonable token budget is allocated (i.e., Figure 1d).
@@
Token Elasticity based Optimal Budget Search. The token elasticity observation shows that while a minimal budget may keep the correctness of the answer, it does not necessarily minimize the token cost. Figure 1c and Figure 1d illustrate an intuitive example. To address this, we enhance Algorithm 1 by incorporating a greedy search strategy aimed at finding the optimal budget that simultaneously minimizes token cost and preserves answer correctness. Specifically, we introduce an additional constraint to the isFeasible condition. Beyond ensuring correctness, the updated budget must result in a lower token cost compared to the previously searched budget. Algorithm 2 outlines the feasibility function employed during the search process. Initially, the actual token cost is computed for both the current and previously evaluated budgets (line 2). Next, feasibility is assessed based on two criteria: the answer correctness and greedy token reduction (line 3). The search process is terminated if either condition fails.
@@ 5.1 Overview
Based on the above analysis, we designed our method TALE for token-budget-aware reasoning in LLMs. Two solutions, i.e., estimation&prompting (TALE-EP, see Figure 4) and post-training (TALE-PT, see Figure 6), are proposed.
Figure 4: The workflow of TALE-EP. Given a question, TALE-EP first estimates the token budget using a budget estimator. It then crafts a token-budget-aware prompt by combining the question with the estimated budget. Finally, the prompt is input to the LLM to generate the answer as the final output. By default, we use the reasoning LLM itself with zero-shot estimation prompt as the budget estimator.
Figure 5: The prompt for zero-shot budget estimation.
5.2 Estimation and Prompting (TALE-EP)
Our observations on token elasticity (Section 4) indicate that only a well-chosen budget within a reasonable range can effectively minimize token costs while preserving LLM performance. The optimal budget, found using Algorithm 1 and Algorithm 2, lies within this range and achieves a satisfying trade-off between efficiency and performance. Building on this insight, we introduce a token budget aware reasoning method by zero-shot-based token budget estimation and prompting the reasoning LLM. TALE-EP leverages the reasoning capabilities of the LLM as an estimator. Figure 4 provides an overview of TALE-EP’s workflow. The goal of TALE-EP is to construct a token-budget-aware prompt that maintains performance comparable to vanilla CoT while reducing token costs. To achieve this balance, TALE-EP follows a two-phase approach: budget estimation and prompt construction. Given a question, TALE-EP first estimates a reasonable token budget that closely aligns with the optimal searched budget. By default, we use the reasoning LLM itself with a zero-shot estimation prompt as the budget estimator. Figure 5 demonstrates the budget estimation prompt. Using this estimate, it then crafts a token-budget-aware prompt and feeds it into the LLM to generate the final answer. Figure 7c illustrates this process with a concrete example. The key intuition behind TALE-EP is inspired by human-like thinking. When solving a mathematical problem, a person may take time to compute the exact answer but can quickly estimate the effort required to solve it. For instance, when comparing a primary school arithmetic question to a college-level calculus problem, one may not immediately provide the solutions but can easily infer that the former takes only seconds while the latter requires significantly more time. Section A.2 evaluates the effectiveness of our budget estimation approach, demonstrating that the budgets estimated by advanced LLMs (e.g., GPT-4o-mini) are generally close to the optimal searched budget and deliver competitive performance.
Figure 6: The workflow of TALE-PT. Given a set of questions, TALE-PT first generates target outputs in phase (1), searching the answers under optimal budget. In phase (2), TALE-PT uses the searched target outputs to craft a specialized dataset, then post-train the LLM (via SFT/DPO) to internalize token-budget awareness.
5.3 TALE Post-Training (TALE-PT)
@@
Another approach for obtaining an LLM with token-budget awareness is post-training it to incorporate this awareness into its inference process, enabling it to generate more token-efficient reasoning responses. Specifically, we post-train the LLM $\mathcal{M}_{\theta}$ to produce answers that adhere to the token budget. This process is divided into two key stages: target output generation and LLM post-training.
Target Output Generation. In the target output generation stage, we craft the target output $y_{i}$ by prompting $\mathcal{M}_{\theta}$ with a Chain-of-Thought (CoT) prompt that incorporates our searched optimal token budget. The prompt is formatted as follows:
@@
where $\beta_{i}^{*}$ is the searched optimal budget for the given question ${\bm{x}}_{i}$ (see search process in Algorithm 1 and Algorithm 2). Figure 1d illustrates an example. The resulting LLM output, constrained by the token budget specified in the prompt, is taken as the crafted target output $y_{i}$. This target output not only leads to the correct answer but also has minimal actual output token cost among our token elasticity-based search process, as described in Section 4. In the LLM post-training stage, we train the LLM $\mathcal{M}_{\theta}$ using the crafted target outputs from the first stage. We introduce two ways to conduct the token-budget awareness internalization during post-training, i.e., SFT-based and DPO-based method. Details of the hype parameters are in Section A.3.
SFT-based Internalization. To inject token-budget awareness into $\mathcal{M}_{\theta}$, we perform supervised fine-tuning with these target outputs. We post-train $\mathcal{M}_{\theta}$ to generate token-efficient outputs by minimizing the cross-entropy loss between the model’s predictions and the target outputs. Given an input ${\bm{x}}$ and a target output $y$ from the first stage (which reflects token-budget awareness), the cross-entropy loss is defined as:
$\mathcal{L}_{\text{CE}}(\theta)=-\frac{1}{N}\sum_{i=1}^{N}\sum_{t=1}^{T_{i}}\log\mathbb{P}(y_{i,t}|y_{i,<t},{\bm{x}}_{i}),$
where $T_{i}$ means the length of the target sequence $y_{i}$ for the $i$-th training example, $y_{i,t}$ the target token at position $t$ of $y_{i}$, $y_{i,<t}$ means the sequence of tokens preceding the current token $y_{i,t}$, representing the context up to time step $t$ for the $i$-th sample. $\mathbb{P}(y_{i,t}|y_{i,<t},x_{i})$ represents the conditional probability predicted by the model $\mathcal{M}_{\theta}$ for the token $y_{i,t}$, given the input ${\bm{x}}_{i}$ and the preceding tokens $y_{i,<t}$. The loss is based on the next token prediction. The goal is to adjust the model parameters $\theta$ such that it produces concise and accurate responses that adhere to the token budget constraint. This is achieved through gradient descent, forcing the model to internalize the compact reasoning patterns from the token-efficient target outputs.
@@
Table 2: Comparison of TALE-EP (estimation and prompting) and other prompt engineering methods. “Directly Answering” means prompting LLM without any reasoning process. “Vanilla CoT” means the vanilla CoT prompting without budget. The model used in our evaluation is GPT-4o-mini OpenAI (2024a). Observe that TALE-EP achieves an average accuracy (ACC) of 80.22%, with an average output token cost of 138.53 and an average expense of 118.46. TALE-EP reduces output token costs by 67%, lowers expenses by 59%, and maintains competitive performance compared to the vanilla CoT approach. ACC $\uparrow$, Output Tokens $\downarrow$, Expense ($10^{-5}\$$ / sample) $\downarrow$.
Dataset Directly Answering Vanilla CoT TALE-EP
@@
Table 3: The generalization of TALE-EP (estimation and prompting) across different LLMs. Yi-lightning Wake et al. (2024), GPT-4o-mini OpenAI (2024a), GPT-4o OpenAI (2024b) and o3-mini OpenAI (2025) are taken into consideration. We conduct following evaluations on the MathBench-College dataset. ACC $\uparrow$, Output Tokens $\downarrow$, Expense ($10^{-5}\$$ / sample) $\downarrow$.
LLM Directly Answering Vanilla CoT TALE-EP